This Question: 1 pt 38 of 60 (37 complete) This Test: 60 pts possib The table provides information about a flu vaccination program. What number of shots per day would a private flu vaccination program provide?f the marginal external benefit from a flu vaccination is equivalent to $15, what is the efficient number of shots? Quantity (shots per day) Marginal cost Marginal benefit (dollars per shot) 100 200 300 400 500 20 30 40 50 60 35 30 25 20 15 A private flu vaccination program would provideshots a day If the marginal external benefit from a flu vaccination is equivalent to $15, the efficient number of shots isa day
Answer #1

A private flu vaccination program would be provided based on market demand and supply, then the number of shots will be provided where marginal private cost equals marginal private benefit.

According to the given marginal private cost equals marginal private benefit at 200 shots where marginal private cost equals marginal private benefits that is $30.

Thus a private flu vaccination program will provide 200 shots per day.

To find out the efficient number of shots per day, then we find the quantity of shots at which marginal social benefit equals marginal cost.

Marginal social benefits = marginal private benefit + external benefit

Quantity Marginal cost Marginal benefit External benefit Marginal social benefit
0 0 0 15 0 + 15 = 15
100 20 35 15 35 + 15 = 50
200 30 30 15 30 + 15 = 45
300 40 25 15 25 + 15 = 40
400 50 20 15 35
500 60 15 15 30

From this table we can easily see that MC and Social benefit get equals at 300 shots where marginal cost and  social benefit equals $40.

So, if the marginal external benefit from a flu vaccination is equivalent to $15,the efficient number of shots is 300 a day.